A laser pulse with wavelength 520 nm contains 4.85 mJ of energy. How many Photons are in the laser pulse?
1
Expert's answer
2020-09-14T08:18:42-0400

Wave length of laser pulse=520nm=520*10^-9 m

We know the energy of a photon(E)=h"\upsilon" h=plank's constant=6.626*10^-34J

"\upsilon" = frequency of photon=c/"\lambda"

c=velocity of light=3*10^8 m/s

"\lambda" =520*10^-9m

Now energy of photons(E)=h(c/"\lambda" )

Putting all the values in above equation we get,E=3.82*10^-19 J

No of photons in laser pulse=total energy of photons/energy of a photon

=4.85*10^-3/3.82*10^-19=1.27*10^16

hence 1.27*10^16 number of photons present